OpenStudy (anonymous):

Find cos theta and sin theta, theta in (0, pi/2) for which cos 2theta = -1/9

OpenStudy (anonymous):

Do we have a domain to work with?

OpenStudy (anonymous):

"theta in (0, pi/2)"

OpenStudy (anonymous):

got ya

OpenStudy (anonymous):

\[\cos (x)=\sqrt{\frac{8}{18}}\] ?

OpenStudy (anonymous):

Duckking the sign this time? :-)

OpenStudy (anonymous):

Yes, that's right...

OpenStudy (anonymous):

How about sin?

OpenStudy (anonymous):

\[\sqrt{\frac{10}{18}}\]